The Purchase of a car requires a $25,000 loan to be repaid in monthly installments for four years at 12% interest compounded monthly and the general inflation is 6% compounded monthly.
a) Find the actual & constant dollar value of the 20th payment.
b) The total loan payback amount in constant & actual dollars
.
Answer #1

Actual 20th payment=25000*(12%/12)/(1-1/(1+12%/12)^(12*4))=658.3458858

Constant 20th payment=658.3458858/(1+6%/12)^20=595.8444394

Total loan payback in actual dollars=658.3458858*12*4=31600.60252

Total loan payback in constant dollars=658.3458858/(6%/12)*(1-1/(1+6%/12)^(12*4))=28032.57703